# Context-Enhanced LLM-Based Framework for Automatic Test Refactoring

### v1

@@ Abstract.
Test Smells, Refactoring, Large Language Models
@@ 1. Introduction
During the software testing process, test code often suffers from test smells, which originate from a lack of sufficient domain knowledge by software engineers and the adoption of poor design practices when writing test code. Recent studies (Damasceno et al., 2022; Pizzini et al., 2023) have shown that developers tend to prioritize production code over test code, which further contributes to the decline in test code quality. From a software maintenance perspective, test smells in a software system can complicate the test code, making it harder to read, understand, and update. This complexity can ultimately lead to a decline in software product quality and reduced developer productivity. Test smells can be removed through test refactoring—a process of improving the internal structure of the code without altering the software’s external behavior (de Paula and Bonifácio, 2022; Soares et al., 2020; Pizzini et al., 2023). Numerous empirical studies (Soares et al., 2020; Pizzini et al., 2023; Pizzini, 2022; Nagy and Abdalkareem, 2022; Damasceno et al., 2022; Soares et al., 2022; Kashiwa et al., 2021) have highlighted the importance of test refactoring, and developers widely accept the refactored tests, and both developers and testers are increasingly recognizing the negative impact of test smells. They widely agree that test quality improves significantly once test smells are eliminated. Additionally, many studies (Peruma et al., 2020; Yang et al., 2024; Aljedaani et al., 2021; Palomba et al., 2018; Pontillo et al., 2024; Wang et al., 2021; Danphitsanuphan and Suwantada, 2012; Fernandes et al., 2022; Bodea, 2022; Virgínio et al., 2020) have proposed test smell detection tools across programming languages such as Java, Scala, and Python, while also evaluating the impact of test smells on the software development process.
@@
Given the remarkable capabilities of large language models (LLMs) in understanding, generating, and reviewing code, they might have the potential to play a key role in test refactoring tasks. In this paper, we explore the potential of open-source LLMs to automate unit test refactoring and propose a approach for automatically eliminating test smells in software projects. However, several challenges must be addressed to achieve this goal:
@@
In this paper, we propose a context-enhanced, LLM-based automatic test refactoring approach for Java projects. Our proposed approach, named UTRefactor, first extracts the project’s test code and related context information, such as the functions and classes under test. We then build an external knowledge base that supports test smell elimination, which includes test smell definitions and descriptions, DSL-based refactoring rules, and other relevant contextual information. Next, we simulate the manual refactoring process typically followed by developers or testers. Using a chain-of-thought approach, we guide the LLM through understanding the test’s intent, identifying the test smells, and following the DSL-defined refactoring steps to refactor the test code. Additionally, we design a checkpoint mechanism to ensure more thorough smell elimination, particularly when multiple test smells are present in a single test.
@@ 2. Motivation
Developers often prioritize maintaining production code while neglecting the maintenance of test code, which can lead to increased maintenance costs, such as those associated with regression testing. (Pizzini et al., 2023). To illustrate the quality issues in test code, we analyze the Gson project (gso, [n. d.]), a popular Java library developed by Google, which has over 23.2k stars on GitHub and continues to be actively maintained, including its test code. Figure 1 presents a intuitive example from the Gson’s JsonObjectTest class. According to the test smell types defined by tsDetect (tsd, [n. d.]), this test has been identified with three distinct smell: ❶ Eager Test: This is a common and challenging type of smell to refactor. It occurs when a test method invokes several methods from the production code, making it harder to understand, maintain, and modify the test code effectively, as the test’s purpose becomes less clear. In this example, this test suffers from the Eager Test smell because it is testing multiple behaviors—adding(add), removing(remove), and checking(has and get) properties—within a single test method. This violates the principle of Single Responsibility Principle (SRP) (srp, [n. d.]). Each test method should focus on testing a specific behavior or function. Otherwise, it becomes harder to pinpoint the root cause of failures and reducing test clarity. To address this, each of these behaviors should be split into separate test methods to ensure that the test remains focused and maintainable. ❷ Duplicate Assert: In this example, the repeated use of the same parameter propertyName across multiple assertions leads to the identification of a Duplicate Assert smell. Eliminating this smell can be achieved by either splitting the original test or transforming it into a Parameterized Test, which removes the need for duplicate assertions by declaring parameterized values in an annotation and executing the test method multiple times. ❸ Assertion Roulette: This is another common smell, indicating that multiple assertions in a test which are lack of clear identification of the reasons for failure, and make it difficult to locate the failing assertion. To address this, each assertion should be supplemented with descriptive information that clarifies its purpose. This allows developers to quickly identify and fix issues when the test fails.
@@
This section presents the details of our proposed approach UTRefactor. As shown in Figure 2, which can be divided into three main steps:
StepPreprocessing. We begin by extracting the test code from the project and detecting all test smells present. For tests that need to be refactored due to the presence of smells, we also extract relevant code context (such as the tested functions and classes) during this step, treating this context as external knowledge for the LLM.
@@ 3.1. Preprocessing
As shown in Figure 2, during this step, we primarily focus on extracting the test code, detecting the test smells within it, and collecting relevant test context information.
@@ 3.1.1. Test Extraction
First, UTRefactor automatically gathers all test files from the test directory (e.g., src/test) within a given Java project. To avoid unnecessary detection of code smells and refactoring overhead, during the test extraction process, we analyze whether the test files contain at least one method annotated with @Test, and automatically filter out files that do not have the @Test annotation. Besides, to analyze the focal methods in a fine-grained approach and meet the requirements of test smell detection, we collect the corresponding classes under test in this step. Specifically, we strip the prefix and suffix Test from the test file names (e.g., ParserTestParser), and use this as an index to search for matching class files in the src/main directory. If a match is successful, the test file and the corresponding class under test are paired and collected together.
@@ 3.1.2. Test Smell Detection
To detect test smells present in the project, we integrate the tsDetect (Peruma et al., 2020), an automated test smell detection tool for Java software systems.
@@ 3.1.3. Test Context Collection
Since refactoring operations are aimed at optimizing the code structure without altering its functional logic, it is crucial to ensure that the LLM has a sufficient understanding of the test code’s functionality before refactoring. Moreover, Yuan et al. (Yuan et al., 2024) has demonstrated that providing additional test context can significantly enhance the performance of LLMs in test code generation tasks. Inspired by these findings and applying them to the task of refactoring test code, we extract relevant contextual information from tests that exhibit smells and use it as external knowledge to help the LLM better comprehend the original test’s intent, as shown in Table 1.
In this step, we extract the contextual information for each test that requires refactoring. During the subsequent refactoring steps, this information is integrated into the refactoring prompt template as a knowledge source to assist the LLM in accurately understanding the test’s original purpose.
@@ 3.2.1. Test Smell Knowledge
Figure 3. An example of test smell explanation of the LLM.
To mitigate the impact of these inconsistencies on the refactoring process, we provide the LLM with a clear and standardized set of test smell definitions before refactoring. Given that Peruma et al. (Peruma et al., 2020) provide comprehensive definitions, explanations, and code examples for 19 test smells in Java, we use these to build part of the external knowledge base for test smells, providing it to the LLM to eliminate inconsistencies in smell interpretation that could impact the refactoring process. For each smell type, we provide a structured definition, including a description of the test smell, its impact, and a pseudocode example to illustrate the concept. During the refactoring process, these standardized smell definitions guide the LLM’s interpretation of the smells present in the test, ensuring consistent and accurate refactoring results.
@@ 3.2.2. Test Refactoring Rule DSL
To eliminate the smells present in tests, we design corresponding refactoring rules for each type of smell. We draw on previous manual approaches to eliminating test smells and empirical research on test smell refactoring (Alfadel et al., 2023). Based on these insights and best testing practices, we formalize DSL rules for each type of test smell, which form the foundation of the entire refactoring process. In designing these DSLs, we exclude simple smell types such as Default Test, which refers to default test classes automatically created by Android Studio when a project is initialized and do not require refactoring. Similarly, we handle Ignored Test (tests marked as ignored) and Empty Test (tests with empty method bodies) by simply removing the unit tests that exhibit these smells. Finally, we define 13 function-level DSLs for Java test smells, aimed at eliminating smells in unit tests. Since these DSLs determine the methods and steps for eliminating test smells, which is crucial for unit test refactoring, we evaluate the correctness of the DSL design by assessing whether the refactored tests successfully eliminate the smells without introducing new issues. This will be discussed in detail in subsequent sections.
@@
These rules are then provided to the LLM as external knowledge during the refactoring process, guiding the LLM to refactor the test according to the predefined rules.
@@ 3.2.3. Sorting Refactoring Rules
Considering that a single test often exhibit multiple types of smells, it is crucial to determine the appropriate order of refactoring operations, as the sequence can significantly impact the effectiveness of the final test code refactoring. To address this, we categorize the test smell refactoring rules and assign them an execution priority. The LLM is then instructed to eliminate test smells according to their priority, from highest to lowest.
@@
Structural Optimizations take precedence over Functional Optimization because a well-structured test provides a solid foundation for functional improvements. For instance, the Assertion Roulette refactoring rule involves adding descriptive messages to assertions. and the Magic Number Test refactoring rule extracts numeric literals into named variables. However, the Duplicate Assertion refactoring rule eliminates duplicate assertions using parameterized tests, which could render the optimizations of Assertion Roulette unnecessary if performed afterward. Therefore, it is crucial to refactor Duplicate Assertion before applying other functional optimizations.
By following this prioritized refactoring approach, we ensure that the test code is systematically improved, starting with essential removals, followed by structural enhancements, and concluding with functional refinements. This strategy optimizes the LLM’s ability to perform accurate and comprehensive refactoring.
@@ 3.3.2. Checkpoint Mechanism
We observed that when a test contains multiple smells, even with sufficient information provided, the LLM may still miss eliminating some of the smells during refactoring. For example, in the testSizeWhenHasInternal function from the Jsoup project, the test has four identified smells: Assertion Roulette, Eager Test, Duplicate Assert, and Magic Number Test. Despite explicitly instructing the LLM to address all four types of smells in the prompt, in practice, the LLM often fails to fully refactor tests with multiple smells.
To address this issue and mitigate incomplete refactoring by the LLM, we design a checkpoint mechanism that allows the LLM to self-verify the completeness of its refactoring process. Specifically, for each smell in the test, we set up a checkpoint that asks whether the refactored test still exhibits the given issue. If the issue persists, the LLM continues refining the refactoring to eliminate the smell. For tests with multiple smells, we construct a chain of smell checkpoints, as illustrated in Figure 2. In the case of the test containing four smells, we collect the corresponding four checkpoints and integrate them into the prompt. The LLM uses this list to systematically check whether the smells has been fully removed. If any issues remain, the LLM will further adjust its refactoring until the test is completely free of the identified smells. This checkpoint mechanism ensures that even in complex cases with multiple smells, the LLM can perform thorough and effective refactoring, ultimately leading to higher-quality test code.
@@ 3.3.3. Refactoring Test Generation
It is essential for developers or testers to understand which tests have been refactored and the reasons behind these changes. To address this, we generate a detailed test refactoring report in this step. This report includes a list of refactored tests, specifying the file locations and test method names. It also outlines the detected smells for each test and the corresponding refactoring methods applied. Additionally, to facilitate subsequent testing and compilation, we organize the refactored test files and functions according to the original project’s test structure.
@@ 4.2. RQ1: How is the quality of the test code after the refactoring of UTRefactor?
Since the test refactoring process must eliminate test smells without altering the original test functionality, we first verify the quality of the refactored test code to ensure effective refactoring. This includes checking for code errors and ensuring that the functionality remains unchanged. Specifically, we first assess the syntax correctness of the refactored test code. To achieve this, we calculate the compilation pass rate and test execution pass rate of all refactored tests, evaluating whether UTRefactor introduce any errors during the structural changes that lead to compilation or runtime failures. Next, to check whether the functionality of the refactored test code remained intact, we compare the line and branch coverage before and after refactoring for all tests that passed execution. If there are no changes in coverage, this indicate that the refactoring only adjust the code structure without modifying the underlying functionality.
@@
We analyze the tests where UTRefactor introduces errors after refactoring. In the Gson project, testStrictComments contains three assertions (assertStrictError) without descriptive information, resulting in the detection of a AR (Assertion Roulette) smell. During the refactoring process, UTRefactor generates descriptive messages for each of the three assertions. However, a parameter mismatch error occurs during compilation. Upon further analysis, we find that assertStrictError is a built-in assertion in Gson, different from the standard JUnit assertions. This assertion only accepts two parameters: an exception describing the expected behavior and an expectedLocation parameter. It does not support an additional message parameter. This conflicts with UTRefactor’s DSL rule for refactoring AR smells, which requires assertions to include a message parameter, leading to a compilation error and causing the smell to remain.

### v2

# Automated Unit Test Refactoring

@@ Abstract.
Test Smells, Test Refactoring, Large Language Models
copyright: acmlicenseddoi: 10.1145/3715750journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE033publicationmonth: 7ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
During the software testing process, test code often suffers from test smells, which originates from a lack of sufficient domain knowledge by software engineers and the adoption of poor design practices when writing test code. Recent studies (Damasceno et al., 2022; Pizzini et al., 2023) have shown that developers tend to prioritize production code over test code, which further contributes to the decline in test code quality. From a software maintenance perspective, test smells in a software system can complicate the test code, making it harder to read, understand, and update. This complexity can ultimately lead to a decline in software product quality and reduced developer productivity. Test smells can be removed through test refactoring—a process of improving the internal structure of the code without altering the software’s external behavior (de Paula and Bonifácio, 2022; Soares et al., 2020; Pizzini et al., 2023). Numerous empirical studies (Soares et al., 2020; Pizzini et al., 2023; Pizzini, 2022; Nagy and Abdalkareem, 2022; Damasceno et al., 2022; Soares et al., 2022; Kashiwa et al., 2021) have highlighted the importance of test refactoring, and developers widely accept the refactored tests, and both developers and testers are increasingly recognizing the negative impact of test smells. They widely agree that test quality improves significantly once test smells are eliminated. Additionally, many studies (Peruma et al., 2020; Yang et al., 2024; Aljedaani et al., 2021; Palomba et al., 2018; Pontillo et al., 2024; Wang et al., 2021; Danphitsanuphan and Suwantada, 2012; Fernandes et al., 2022; Bodea, 2022; Virgínio et al., 2020) have proposed test smell detection tools across programming languages such as Java, Scala, and Python, while also evaluating the impact of test smells on the software development process.
@@
Given the remarkable capabilities of large language models (LLMs) in understanding, generating, and reviewing code, they might have the potential to play a key role in test refactoring tasks. In this paper, we explore the potential of open-source LLMs to automate unit test refactoring and propose an approach for automatically eliminating test smells in software projects. However, several challenges must be addressed to achieve this goal:
@@
In this paper, we propose a context-enhanced, LLM-based automatic test refactoring approach for Java projects. Our proposed approach, named UTRefactor, first extracts the project’s test code and related context information, such as the methods and classes under test. We then build an external knowledge base that supports test smell elimination, which includes test smell definitions and descriptions, DSL-based refactoring rules, and other relevant contextual information. Next, we simulate the manual refactoring process typically followed by developers or testers. Using a chain-of-thought approach, we guide the LLM through understanding the test’s intent, identifying the test smells, and following the DSL-defined refactoring steps to refactor the test code. Additionally, we design a checkpoint mechanism to ensure more thorough smell elimination, particularly when multiple test smells are present in a single test.
@@ 2. Motivation
Developers often prioritize maintaining production code while neglecting the maintenance of test code, which can lead to increased maintenance costs, such as those associated with regression testing. (Pizzini et al., 2023). To illustrate the quality issues in test code, we analyze the Gson project (gso, [n. d.]), a popular Java library developed by Google, which has over 23.2k stars on GitHub and continues to be actively maintained, including its test code. Figure 1 presents an intuitive example from the Gson’s JsonObjectTest class. According to the test smell types defined by tsDetect (tsd, [n. d.]), this test has been identified with three distinct smells:
Eager Test: This is a common and challenging type of smell to refactor. It occurs when a test method invokes several methods from the production code, making it harder to understand, maintain, and modify the test code effectively, as the test’s purpose becomes less clear. In this example, this test suffers from the Eager Test smell because it is testing multiple behaviors—adding(add), removing(remove), and checking(has and get) properties—within a single test method. This violates the principle of Single Responsibility Principle (SRP) (srp, [n. d.]). Each test method should focus on testing a specific behavior or function. Otherwise, it becomes harder to pinpoint the root cause of failures and reduce test clarity. To address this, each of these behaviors should be split into separate test methods to ensure that the test remains focused and maintainable.
@@
Duplicate Assert: In this example, the repeated use of the same parameter propertyName across multiple assertions leads to the identification of a Duplicate Assert smell. Eliminating this smell can be achieved by either splitting the original test or transforming it into a Parameterized Test, which removes the need for duplicate assertions by declaring parameterized values in an annotation and executing the test method multiple times.
Assertion Roulette: This is another common smell, indicating that multiple assertions in a test lack clear identification of the reasons for failure, and make it difficult to locate the failing assertion. To address this, each assertion should be supplemented with descriptive information that clarifies its purpose. This allows developers to quickly identify and fix issues when the test fails.
@@
This section presents the details of our proposed approach UTRefactor. As shown in Figure 2, it can be divided into three main steps:
StepPreprocessing. We begin by extracting the test code from the project and detecting all test smells present. For tests that need to be refactored due to the presence of smells, we also extract relevant code context (such as the tested methods and classes) during this step, treating this context as external knowledge for the LLM.
@@ 3.1. Preprocessing
As shown in Figure 2, this step involves three key processes: extracting the test code, identifying test smells, and gathering relevant contextual information about the tests.
@@ 3.1.1. Test Extraction
First, UTRefactor automatically gathers all test files from the test directory (e.g., src/test) within a given Java project. To avoid unnecessary detection of code smells and refactoring overhead, during the extraction process, we analyze whether the test files contain at least one method annotated with @Test and automatically filter out files that do not have the @Test annotation. Besides, to analyze the focal methods in a fine-grained approach and meet the requirements of test smell detection, we collect the corresponding classes under test in this step. Specifically, we strip the prefix and suffix Test from the test file names (e.g., ParserTestParser), and use this as an index to search for matching class files in the src/main directory. If a match is successful, the test file and the corresponding class under test are paired and collected together.
@@ 3.1.2. Test Smell Detection
To detect test smells present in the project, we integrate the tsDetect (Peruma et al., 2020), an automated test smell detection tool for Java projects.
@@ 3.1.3. Test Context Collection
Since refactoring operations are aimed at optimizing the code structure without altering its functional logic, it is crucial to ensure that the LLM has a sufficient understanding of the test code’s functionality before refactoring. Moreover, Yuan et al. (Yuan et al., 2024) have demonstrated that providing additional test context can significantly enhance the performance of LLMs in test code generation tasks. Inspired by these findings and applying them to the task of refactoring test code, we extract relevant contextual information from tests that exhibit smells and use it as external knowledge to help the LLM better comprehend the original test’s intent, as shown in Table 1.
In this step, we extract the contextual information for each test that requires refactoring. During the subsequent refactoring steps, this information is integrated into the refactoring prompt template as a knowledge source to assist the LLM in accurately understanding the test’s original intent.
@@ 3.2.1. Test Smell Knowledge
To mitigate the impact of these inconsistencies on the refactoring process, we provide the LLM with a clear and standardized set of test smell definitions before refactoring. Given that Peruma et al. (Peruma et al., 2020) provide comprehensive definitions, explanations, and code examples for 19 test smells in Java, we use these to build part of the external knowledge base for test smells, providing it to the LLM to eliminate inconsistencies in smell interpretation that could impact the refactoring process. Due to space limitations, the detailed definitions of these smells are available in our replication package (utr, [n. d.]). For each smell type, we provide a structured definition, including a description of the test smell, its impact, and a pseudocode example to illustrate the concept. During the refactoring process, these standardized smell definitions guide the LLM’s interpretation of the smells present in the test, ensuring consistent and accurate refactoring results.
@@ 3.2.2. Test Refactoring Rule DSL
To eliminate the smells present in tests, we design corresponding refactoring rules for each type of smell. We draw on previous manual approaches to eliminating test smells and empirical research on test smell refactoring (Alfadel et al., 2023). Based on these insights and best testing practices, we formalize DSL rules for each type of test smell, which form the foundation of the entire refactoring process. In designing these DSLs, we exclude simple smell types such as Default Test, which refers to default test classes automatically created by Android Studio when a project is initialized and do not require refactoring. Similarly, we handle Ignored Test (tests marked as ignored) and Empty Test (tests with empty method bodies) by simply removing the unit tests that exhibit these smells. Finally, we define 13 method-level DSLs for Java test smells, aimed at eliminating smells in unit tests. Due to space limitations, the complete set of DSLs is available in our replication package (utr, [n. d.]). Since these DSLs determine the methods and steps for eliminating test smells, which is crucial for unit test refactoring, we evaluate the correctness of the DSL design by assessing whether the refactored tests successfully eliminate the smells without introducing new issues. This will be discussed in detail in subsequent sections.
@@
We put the details of DSLs on our replication package (utr, [n. d.]). These rules are then provided to the LLM as external knowledge during the refactoring process, guiding the LLM to refactor the test according to the predefined rules.
@@ 3.2.3. Sorting Refactoring Rules
Considering that a single test often exhibits multiple types of smells, it is crucial to determine the appropriate order of refactoring operations, as the sequence can significantly impact the effectiveness of the final test code refactoring. To address this, we categorize the test smell refactoring rules and assign them an execution priority. The LLM is then instructed to eliminate test smells according to their priority, from highest to lowest.
@@
Structural Optimizations take precedence over Functional Optimization because a well-structured test provides a solid foundation for functional improvements. For example, the Assertion Roulette refactoring rule involves adding descriptive messages to assertions, and the Magic Number Test refactoring rule extracts numeric literals into named variables. However, the Duplicate Assertion refactoring rule eliminates duplicate assertions using parameterized tests, which could render the optimizations of Assertion Roulette unnecessary if performed afterward. Therefore, it is crucial to refactor Duplicate Assertion before applying other functional optimizations.
By following this prioritized refactoring approach, we ensure that the test code is improved in an orderly manner, starting with essential removals, followed by structural enhancements, and concluding with functional refinements.
@@ 3.3.2. Checkpoint Mechanism
When a test contains multiple smells, even with sufficient information provided, the LLM may still miss eliminating some of the smells during refactoring. For example, in the testSizeWhenHasInternal method from the Jsoup project, the test has four identified smells: Assertion Roulette, Eager Test, Duplicate Assert, and Magic Number Test. Despite explicitly instructing the LLM to address all four types of smells in the prompt, in practice, the LLM often fails to fully refactor tests with multiple smells.
To address this issue and mitigate incomplete refactoring by the LLM, we design a checkpoint mechanism that allows the LLM to self-verify the completeness of its refactoring process. Specifically, for each smell in the test, we set up a checkpoint that asks whether the refactored test still exhibits the given issue. If the issue persists, the LLM continues refining the refactoring to eliminate the smell. For tests with multiple smells, we construct a chain of smell checkpoints, as illustrated in Figure 2. In the case of the test containing four smells, we collect the corresponding four checkpoints and integrate them into the prompt. The LLM uses this list to systematically check whether the smells have been fully removed. If any issues remain, the LLM will further adjust its refactoring until the test is completely free of the identified smells. This checkpoint mechanism ensures that even in complex cases with multiple smells, the LLM can perform thorough and effective refactoring, ultimately leading to higher-quality test code.
@@ 3.3.3. Refactoring Test Generation
It is essential for developers or testers to understand which tests have been refactored and the reasons behind these changes. To address this, we generate a detailed test refactoring report in this step. This report includes a list of refactored tests, specifying the file locations and test method names. It also outlines the detected smells for each test and the corresponding refactoring methods applied. Additionally, to facilitate subsequent testing and compilation, we organize the refactored test files and methods according to the original project’s test structure.
@@ 4.2. RQ1: How is the quality of the test code after the refactoring of UTRefactor?
Since the test refactoring process must eliminate test smells without altering the original test functionality, we first verify the quality of the refactored test code to ensure effective refactoring. This includes checking for code errors and ensuring that the functionality remains unchanged. Specifically, we first assess the syntax correctness of the refactored test code. To achieve this, we calculate the compilation pass rate and test execution pass rate of all refactored tests, evaluating whether UTRefactor introduces any errors during the structural changes that lead to compilation or runtime failures.
@@
We analyze the tests where UTRefactor introduces errors after refactoring. In the Gson project, testStrictComments contains three assertions (assertStrictError) without descriptive information, resulting in the detection of an AR (Assertion Roulette) smell. During the refactoring process, UTRefactor generates descriptive messages for each of the three assertions. However, a parameter mismatch error occurs during compilation. Upon further analysis, we find that assertStrictError is a built-in assertion in Gson, different from the standard JUnit assertions. This assertion only accepts two parameters: an exception describing the expected behavior and an expectedLocation parameter. It does not support an additional message parameter. This conflicts with UTRefactor’s DSL rule for refactoring AR smells, which requires assertions to include a message parameter, leading to a compilation error and causing the smell to remain.
@@
Consistency of Refactoring. Given the challenge of determining functional and behavioral consistency of test code before and after refactoring, we employ a more advanced model (GPT-4o) to evaluate the functional consistency of test code pre- and post-refactoring, supplemented by a manual secondary review of the evaluation results. Specifically, we provide GPT-4o with all 879 pairs of test code that can be executed successfully before and after refactoring to assess whether the changes involve only structural adjustments while maintaining functional consistency.
Table 4. Classification of test inconsistencies before and after refactoring.
Inconsistent Type Num
Exception Handling 18
Redundant Assertion 9
Redundant Variable Declaration 5
Table 4 indicates that 32 cases (approximately 4%) are flagged as Inconsistent, which can be categorized into three types: Firstly, regarding differences in exception handling (18 cases), although GPT-4o identifies a behavioral change between the original test code’s throws exception handling and the DSL-refactored JUnit assertDoesNotThrow method, this refactoring actually aligns with the design philosophy of the JUnit 5 framework, providing a more standardized exception verification mechanism.
Secondly, in terms of redundant assertion handling (9 cases), the current DSL strategy of directly deleting redundant assertions is marked by GPT-4o as a potential behavioral change, prompting us to plan an optimization of the refactoring rules to comment out redundant assertions and add developer prompts. Lastly, concerning redundant variable declarations (5 cases), variables defined during the setup phase are redundantly declared in subtests when splitting the original test into multiple subtests. Although there is still room for improvement in UTRefactor’s handling of redundant assertions and duplicate variables, it performs excellently overall in maintaining test functional consistency.